40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 9 13:40:02 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics 
    Reply  Tue Nov 9 16:55:52 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics IMG_8107.JPG
       Reply  Tue Nov 9 18:05:03 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics 
          Reply  Mon Nov 15 15:12:28 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics 
             Reply  Tue Nov 16 17:29:49 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics x_x_TF1.pdf
                Reply  Thu Nov 18 20:00:43 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics Final_Plant_Testing.zip
                   Reply  Mon Nov 22 16:38:26 2021, Tega, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics 
                      Reply  Mon Jan 24 17:10:25 2022, Tega, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics Screen_Shot_2022-01-24_at_5.33.15_PM.png
                         Reply  Thu Jan 27 16:40:57 2022, Tega, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics simplant_control_1.pngsimplant_control_0.png
                            Reply  Thu Feb 3 17:15:23 2022, Tega, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics Screen_Shot_2022-02-03_at_8.23.07_PM.png
                               Reply  Wed Feb 9 14:34:27 2022, Ian, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics SummaryMdemScreen.png
Message ID: 16645     Entry time: Thu Feb 3 17:15:23 2022     In reply to: 16626     Reply to this: 16654
Author: Tega 
Type: Summary 
Category: Computer Scripts / Programs 
Subject: SUS Plant Plan for New Optics 

Finally got the SIMPLANT damping to work following Rana's suggestion to try damping one DoF at a time, woo-hoo!

At first, things didn't look good even when we only focus on the POS DoF. I then noticed that the input value (X1:SUS-OPT_PLANT_TM_RESP_1_1_IN1) to the plant was always zero. This was odd bcos it meant the control signal was not making its way to the plant. So I decided to look at the sensor data

(X1:SUS-OPT_PLANT_COIL_IN_UL_OUTPUT, X1:SUS-OPT_PLANT_COIL_IN_UR_OUTPUT, X1:SUS-OPT_PLANT_COIL_IN_LR_OUTPUT, X1:SUS-OPT_PLANT_COIL_IN_LL_OUTPUT)

that adds up via the C2DOF matrix to give the POS DoF and I noticed that these interior nodes can take on large values but always sum up to zero because the pair (UL, LL) was always the negative of (UR,LR). These things should have the same sign, at least in our case where only the POS DoF is excited, so I tracked the issue back to the alternating (-,+,-,+,-) convention for the gains

(X1:SUS-OPT_CTRL_ULCOIL_GAIN, X1:SUS-OPT_CTRL_URCOIL_GAIN, X1:SUS-OPT_CTRL_LRCOIL_GAIN, X1:SUS-OPT_CTRL_LLCOIL_GAIN, X1:SUS-OPT_CTRL_SDCOIL_GAIN)

of the Coil Output filters used in the real system, which we adopted in the hopes that all was well. Anyways, I changed them all back to +1. This also means that we need to change the sign of the gain for the SIDE filter, which I have done also (and check that it damps OK). I decided to reduce the magnitude of the SIDE damping from 1 to 0.1 so that we can see the residuals since the value of -1 quickly sends the error to zero.  I also increased the gain magnitude for the other DoF to 4.

When looking at the plot remember that the values actually represent counts with a scaling of 2^15 (or 32768) from the ADC. I switched back to the original filters on FM1 (e.g. pit_pit ) without damping coefficients present in the FM2 filter (e.g. pit_pit_damp).


FYI, Rana used the ETMY suspension MEDM screen to illustrate the working of the single suspension to me and changed maybe POS and PITCH gains while doing so.


Also, the Medify purifier 'replace filter' indicator issue occurred bcos the moonlight button should have been pressed for 3 seconds to reset the 'replace filter' indicator after filter replacement.

Attachment 1: Screen_Shot_2022-02-03_at_8.23.07_PM.png  213 kB  Uploaded Thu Feb 3 20:24:58 2022  | Show | Show all
ELOG V3.1.3-